Water flows into a tank at a rate of 15 litres per minute. At the beginning the tank contained 250 litres.
a) Write an expression for the volume, V litres, of water in the tank at time t minutes.
b) How many litres of water are there in the tank after an hour?
c) The tank has a capacity of 5000 litres. How long does it take to fill?

Answer by mananth(16949)

Water in tank = 250 liters
rate of flow = 15 liters/minute
...
volume in tank after time t, V= 250+15t
...
V= 250 +15*60
V=250+900
V= 1150 liters after 1 hour.
...
5000 = 250 +15*t
5000-250=15*t
4750 = 15*t
/15
4750/15 = t
t=316 minutes
...
m.ananth@hotmail.ca